# Genuine and Robust Magnetoelectric Coupling Effect in van der Waals Multiferroic Tunnel Junctions

### v1

@@ Abstract
Van der Waals multiferroic tunnel junctions (vdW-MFTJs) are promising candidates for data storage devices due to their tunable thickness and capability to exhibit multiple nonvolatile resistance states. However, existing vdW-MFTJs utilize ferroelectricity and ferromagnetism independently, failing to achieve true magnetoelectric coupling, which leads to unnecessary energy dissipation. Here, we propose an innovative vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure, enabling genuine magnetoelectric coupling without relying on atomic migration induced by inversion symmetry breaking for ferroelectric polarization reversal. Using first-principles calculations, we investigate the spin-polarized quantum transport properties of the proposed structure. By integrating asymmetric PtTe2/alkali-metal (Li/Na/K)-doped/intercalated CrBr3 electrodes, the device demonstrates exceptional performance, with a maximum tunneling magnetoresistance (TMR) exceeding % and tunneling electroresistance (TER) reaching 2499%, while the spin-filtering channels can be flexibly controlled by the magnetization direction of the magnetic free layer, achieving perfect spin-filtering over a broad bias voltage range. Applying an external bias voltage further enhances these metrics, increasing TMR to % and TER to 9990%. Notably, a pronounced negative differential resistance (NDR) effect is observed, yielding an unprecedented peak-to-valley ratio (PVR) of %, the highest value reported to date. These extraordinary characteristics highlight the potential of vdW-MFTJs for ultra-efficient electronic switching, a key feature for next-generation spintronic devices. Our findings provide a solid theoretical foundation for designing and developing high-performance magnetic storage and logic technologies.
@@
In this Letter, we present a novel vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure, offering genuine magnetoelectric coupling without relying on conventional atomic migration induced by inversion symmetry breaking for polarization reversal. Employing first-principles calculations, we investigated its spin-polarized quantum transport properties and demonstrated remarkable performance enhancements through the integration of asymmetric PtTe2/alkali-metal (Li/Na/K)-doped/intercalated CrBr3 electrodes. The proposed design exhibits exceptional tunneling magnetoresistance (TMR) and tunneling electroresistance (TER), reaching 8.1105% and 2499%, respectively, along with perfect spin-filtering effects over a broad bias voltage range. Furthermore, applying an external bias voltage amplifies these metrics to unprecedented levels, with TMR and TER increasing to 3.6107% and 9990%, respectively. Notably, the device displays a pronounced negative differential resistance (NDR) effect, characterized by a peak-to-valley ratio (PVR) as high as 9.55109%, the largest reported to date. These extraordinary properties underscore the transformative potential of vdW-MFTJs for ultra-efficient electronic switching, laying a robust theoretical foundation for next-generation spintronic devices and advancing high-performance magnetic storage and logic technologies.
@@
Electronic and transport properties at equilibrium.—The mechanism of traditional MFTJs relies on external magnetic or electric fields to switch the polarization direction of the magnetic free layer or electrostatic barrier, resulting in four non-volatile resistance states: the magnetic ordering-dependent parallel (PC) and antiparallel (APC) configurations, and the ferroelectric polarization-dependent P and P states. In contrast, due to the unique magnetoelectric coupling in the CrBr3/MnPSe3/CrBr3 heterostructure, the control of these four resistance states can be achieved solely by an external magnetic field, enabling low-power operation of our MFTJs. We first investigated the electronic transport properties of PtTe2/CrBr3/MnPSe3/CrBr3/CrBr3(Li/Na/K) vdW-MFTJs at equilibrium. As summarized in Table 1, Li/Na/K doping induces significant TMR and TER effects at equilibrium, with maximum values of 8.1105% and 2499%, respectively. Furthermore, all four non-volatile resistance states exhibit perfect spin filtering effects, and the direction of the magnetic free layer can flexibly switch the spin channels. The experimentally feasible Li intercalation strategy for CrBr3 is detailed in Table S1 of the Supplemental Material [38]. Despite a one-order-of-magnitude reduction in TMR/TER, Li intercalation preserves the perfect spin filtering effect, with the TMR/TER still remaining substantial. Therefore, these vdW-MFTJs exhibit giant TMR/TER and perfect spin filtering effects with tunable spin channels, showcasing their significant potential for spintronic applications.
@@
Transport properties at nonequilibrium.—In this section, we calculate the bias voltage-dependent (ranging from to V) spin-polarized current, spin injection efficiency (SIE), TMR, and TER for the MFTJs under four resistance states, as illustrated in Fig. 4. From Fig. 4(a), one can observe that in the P state with PC magnetic alignment, the total current is primarily contributed by the spin-up current, while the spin-down current is significantly suppressed, demonstrating an intriguing spin-filtering effect. The SIE remains 100% across the entire bias voltage range [see Fig. 4(d)]. Moreover, the total current increases with the bias voltage, then sharply decreases when the bias exceeds V ( V), indicating the presence of a typical negative differential resistance (NDR) effect. The current peak-to-valley ratio (PVR) is crucial for applications of NDR devices. For instance, in high-frequency applications, the PVR determines the efficiency of converting d.c.-to-RF signals [39], while in multi-valued logic it defines the voltage range for transitioning to intermediate states [40]. Specifically, the current peaks at V and reaches a minimum at V, with a peak-to-valley difference spanning over ten orders of magnitude, resulting in an exceptionally high PVR of 9.55109% [see Fig. 4(h)], the highest value reported to date. As displayed in Fig. 4(b), in the APC state, the spin-filtering channel is reversed, i.e., spin-up electron transport is blocked, and the NDR effect occurs under negative bias. Comparing Figs. 4(a) and 4(b), it is clear that the current in the PC state is significantly larger than in the APC state, indicating a high TMR. For the P state, the behavior is similar to that of the P state. More intriguingly, the distinction lies in the spin-filtering channel being reversed once again as depicted in Figs. 4(e) and 4(f), i.e., the PC/APC states conduct spin-down/spin-up currents, respectively. This demonstrates that the spin-filtering channel can be flexibly manipulated among the four resistance states. Furthermore, the current in the P state is substantially larger than in the P state for both magnetic configurations, implying the realization of a large TER. From the I-V curves, the bias-dependent evolution of TMR and TER is derived, as shown in Figs. 4(c) and 4(g). The maximum TMR (3.6107%) and TER (9990%) are achieved at and V, respectively, for the P/APC state.
@@
Summary and Discussion.—In summary, we propose a novel vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure that exhibits genuine magnetoelectric coupling, a significant breakthrough in spintronic device engineering. This design overcomes the traditional limitations by enabling efficient spin and charge modulation through polarization switching, without the need for atomic migration leading to symmetry breaking, as in conventional ferroelectric systems. The incorporation of asymmetric PtTe2/alkali metal-doped/intercalated CrBr3 electrodes leads to extraordinary tunneling magnetoresistance (TMR) and tunneling electroresistance (TER) values, exceeding 8.1105% and 2499%, respectively, at zero bias. Under external bias, these values are dramatically enhanced, reaching up to 3.6 107% for TMR and 9990% for TER. Furthermore, the device demonstrates an exceptional negative differential resistance (NDR) effect, with a peak-to-valley ratio (PVR) of 9.55109%, the highest reported to date. These outstanding properties underscore the transformative potential of vdW-MFTJs for energy-efficient electronic switching, as well as high-performance data storage and logic devices.

### v2

@@ Abstract
Van der Waals multiferroic tunnel junctions (vdW-MFTJs) are promising candidates for data storage devices due to their tunable thickness and capability to exhibit multiple nonvolatile resistance states. However, existing vdW-MFTJs utilize ferroelectricity and ferromagnetism independently, failing to achieve true magnetoelectric coupling, which leads to unnecessary energy dissipation. Here, we propose an innovative vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure, enabling genuine magnetoelectric coupling without relying on atomic migration induced by inversion symmetry breaking for ferroelectric polarization reversal. Using first-principles calculations, we investigate the spin-polarized quantum transport properties of the proposed structure. By integrating asymmetric PtTe2/alkali-metal (Li/Na/K)-doped/intercalated CrBr3 electrodes, the device demonstrates exceptional performance, with a maximum tunneling magnetoresistance (TMR) exceeding % and tunneling electroresistance (TER) reaching 2499%, while the spin-filtering channels can be flexibly controlled by the magnetization direction of the magnetic free layer, achieving perfect spin-filtering over a broad bias voltage range. Applying an external bias voltage further enhances these metrics, increasing TMR to % and TER to 9990%. Notably, a pronounced negative differential resistance (NDR) effect is observed, yielding an unprecedented peak-to-valley ratio (PVR) of %, representing the highest value reported for vertical tunnel junctions. These extraordinary characteristics highlight the potential of vdW-MFTJs for ultra-efficient electronic switching, a key feature for next-generation spintronic devices. Our findings provide a solid theoretical foundation for designing and developing high-performance magnetic storage and logic technologies.
@@
In this Letter, we present a novel vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure, offering genuine magnetoelectric coupling without relying on conventional atomic migration induced by inversion symmetry breaking for polarization reversal. Employing first-principles calculations, we investigated its spin-polarized quantum transport properties and demonstrated remarkable performance enhancements through the integration of asymmetric PtTe2/alkali-metal (Li/Na/K)-doped/intercalated CrBr3 electrodes. The proposed design exhibits exceptional tunneling magnetoresistance (TMR) and tunneling electroresistance (TER), reaching 8.1105% and 2499%, respectively, along with perfect spin-filtering effects over a broad bias voltage range. Furthermore, applying an external bias voltage amplifies these metrics to unprecedented levels, with TMR and TER increasing to 3.6107% and 9990%, respectively. Notably, the device displays a pronounced negative differential resistance (NDR) effect, characterized by a peak-to-valley ratio (PVR) as high as 9.55109%, the largest reported in vertical tunnel junctions. These extraordinary properties underscore the transformative potential of vdW-MFTJs for ultra-efficient electronic switching, laying a robust theoretical foundation for next-generation spintronic devices and advancing high-performance magnetic storage and logic technologies.
@@
Electronic and transport properties at equilibrium.—The mechanism of traditional MFTJs relies on external magnetic or electric fields to switch the polarization direction of the magnetic free layer or electrostatic barrier, resulting in four non-volatile resistance states: the magnetic ordering-dependent parallel (PC) and antiparallel (APC) configurations, and the ferroelectric polarization-dependent P and P states. In contrast, due to the unique magnetoelectric coupling in the CrBr3/MnPSe3/CrBr3 heterostructure, the control of these four resistance states can be achieved solely by an external magnetic field, enabling low-power operation of our MFTJs. We first investigated the electronic transport properties of PtTe2/CrBr3/MnPSe3/CrBr3/CrBr3(Li/Na/K) vdW-MFTJs at equilibrium. As summarized in Table 1, Li/Na/K doping induces significant TMR and TER effects at equilibrium, with maximum values of 8.1105% and 2499%, respectively. Furthermore, all four non-volatile resistance states exhibit perfect spin filtering effects, and the direction of the magnetic free layer can flexibly switch the spin channels. The experimentally feasible Li intercalation strategy for CrBr3 is detailed in Table SI of the Supplemental Material [38]. Despite a one-order-of-magnitude reduction in TMR/TER, Li intercalation preserves the perfect spin filtering effect, with the TMR/TER still remaining substantial. Therefore, these vdW-MFTJs exhibit giant TMR/TER and perfect spin filtering effects with tunable spin channels, showcasing their significant potential for spintronic applications.
@@
Transport properties at nonequilibrium.—In this section, we calculate the bias voltage-dependent (ranging from to V) spin-polarized current, spin injection efficiency (SIE), TMR, and TER for the MFTJs under four resistance states, as illustrated in Fig. 4. From Fig. 4(a), one can observe that in the P state with PC magnetic alignment, the total current is primarily contributed by the spin-up current, while the spin-down current is significantly suppressed, demonstrating an intriguing spin-filtering effect. The SIE remains 100% across the entire bias voltage range [see Fig. 4(d)]. Moreover, the total current increases with the bias voltage, then sharply decreases when the bias exceeds V ( V), indicating the presence of a typical negative differential resistance (NDR) effect. The current peak-to-valley ratio (PVR) is crucial for applications of NDR devices. For instance, in high-frequency applications, the PVR determines the efficiency of converting d.c.-to-RF signals [39], while in multi-valued logic it defines the voltage range for transitioning to intermediate states [40]. Specifically, the current peaks at V and reaches a minimum at V, with a peak-to-valley difference spanning over ten orders of magnitude, resulting in an exceptionally high PVR of 9.55109% [see Fig. 4(h)], the highest value reported to date for vertical tunnel junctions. We have also summarized a series of reported PVR values in Table SII in the Supplemental Material [38]. The current highest value, observed in the N-doped monolayer GeS system [41], is two orders of magnitude smaller than ours. As displayed in Fig. 4(b), in the APC state, the spin-filtering channel is reversed, i.e., spin-up electron transport is blocked, and the NDR effect occurs under negative bias. Comparing Figs. 4(a) and 4(b), it is clear that the current in the PC state is significantly larger than in the APC state, indicating a high TMR. For the P state, the behavior is similar to that of the P state. More intriguingly, the distinction lies in the spin-filtering channel being reversed once again as depicted in Figs. 4(e) and 4(f), i.e., the PC/APC states conduct spin-down/spin-up currents, respectively. This demonstrates that the spin-filtering channel can be flexibly manipulated among the four resistance states. Furthermore, the current in the P state is substantially larger than in the P state for both magnetic configurations, implying the realization of a large TER. From the I-V curves, the bias-dependent evolution of TMR and TER is derived, as shown in Figs. 4(c) and 4(g). The maximum TMR (3.6107%) and TER (9990%) are achieved at and V, respectively, for the P/APC state.
@@
Summary and Discussion.—In summary, we propose a novel vdW-MFTJ design based on a CrBr3/MnPSe3/CrBr3 vertical heterostructure that exhibits genuine magnetoelectric coupling, a significant breakthrough in spintronic device engineering. This design overcomes the traditional limitations by enabling efficient spin and charge modulation through polarization switching, without the need for atomic migration leading to symmetry breaking, as in conventional ferroelectric systems. The incorporation of asymmetric PtTe2/alkali metal-doped/intercalated CrBr3 electrodes leads to extraordinary tunneling magnetoresistance (TMR) and tunneling electroresistance (TER) values, exceeding 8.1105% and 2499%, respectively, at zero bias. Under external bias, these values are dramatically enhanced, reaching up to 3.6 107% for TMR and 9990% for TER. Furthermore, the device demonstrates an exceptional negative differential resistance (NDR) effect, with a peak-to-valley ratio (PVR) of 9.55109%, the highest reported to date for vertical tunnel junctions. These outstanding properties underscore the transformative potential of vdW-MFTJs for energy-efficient electronic switching, as well as high-performance data storage and logic devices.